# Sensor Applications

### v1

@@ Abstract
Multispectral object detection aims to leverage complementary information from visible (RGB) and infrared (IR) modalities to enable robust performance under diverse environmental conditions. In this letter, we propose IC-Fusion, a multispectral object detector that effectively fuses visible and infrared features through a lightweight and modality-aware design. Motivated by wavelet analysis and empirical observations, we find that IR images contain structurally rich high-frequency information critical for object localization, while RGB images provide complementary semantic context. To exploit this, we adopt a compact RGB backbone and design a novel fusion module comprising a Multi-Scale Feature Distillation (MSFD) block to enhance RGB features and a three-stage fusion block with Cross-Modal Channel Shuffle Gate (CCSG) and Cross-Modal Large Kernel Gate (CLKG) to facilitate effective cross-modal interaction. Experiments on the FLIR and LLVIP benchmarks demonstrate the effectiveness and efficiency of our IR-centric fusion strategy. Our code is available at https://github.com/smin-hwang/IC-Fusion.
{IEEEkeywords}
@@ 1 Introduction
Despite the potential of multispectral object detection [1, 2, 3, 4, 5], effectively integrating the two modalities remains challenging due to their inherent spectral differences. To better understand the contribution of each modality, we perform a wavelet decomposition on RGB and IR images, as shown in Fig. 1. The RGB image primarily contains low-frequency semantic structures in the LL band, while its high-frequency components are often sparse and noisy. In contrast, the IR modality presents distinct object boundaries and contours, especially in the LH and HL sub-bands, which are crucial for robust localization and shape understanding. This observation highlights that IR images inherently encode sharper structural cues that are highly beneficial for object detection, particularly in edge and boundary regions.
@@
The effectiveness of IR features is further quantitatively validated in Table 1, where unimodal IR detectors consistently outperform their RGB counterparts on the FLIR-aligned [6] benchmark across various backbones. For example, RT-DETR [7] trained solely on IR data achieves a mAP of 43.6, significantly surpassing its RGB equivalent (33.5 mAP). These findings imply that heavy feature extraction from the RGB modality may be redundant and detection performance can be retained or improved, even with significantly reduced computation by relying on IR-centric representations and adopting a lightweight RGB backbone.
Motivated by these insights, we propose a multispectral object detection transformer, IC-Fusion, which incorporates a novel cross-modal fusion module to efficiently fuse information across modalities. To reduce redundancy and improve efficiency, we adopt a lightweight backbone for the RGB modality while preserving a deeper IR representation to retain structural details. Our design emphasizes the dominant structural contributions of the IR modality while enhancing the complementary semantic information from RGB. Specifically, we introduce a Multi-Scale Feature Distillation (MSFD) module to enrich RGB features with multi-scale spatial context, and adopt a three-stage fusion block consisting of two Cross-Modal Channel Shuffle Gates (CCSG) and one Cross-Modal Large Kernel Gate (CLKG) to effectively extract and fuse complementary cross-modal representations. Experiments on benchmark datasets validate the effectiveness of our design and further highlight the advantage of IR-centric fusion for efficient multispectral object detection.
Table 1: Comparisons on the FLIR-aligned dataset. We report mAP@50, mAP@75, and mAP under IoU thresholds.
@@
Figure 2: Overview of (a) the proposed IC-Fusion framework and (b) the proposed cross-modal feature fusion module composed of CCSG and CLKG blocks.
2 Proposed Method
2.1 Overview
We propose IC-Fusion, a multispectral object detection transformer framework that effectively integrates RGB and IR features through structured cross-modal fusion. As shown in Fig. 2, the framework consists of three main components: dual-stream backbones, a fusion module, and a DETR-based transformer encoder-decoder.
@@ 2.1 Overview
where $\mathcal{F}(\cdot)$ represents the cross-modal fusion process, which sequentially applies a Cross-Modal Channel Shuffle Gate (CCSG) for feature interaction and a Cross-Modal Large Kernel Gate (CLKG) for long-range contextual alignment. The fused output integrates complementary semantics from both modalities while suppressing redundant information, and is subsequently passed to an efficient encoder [7], followed by a DETR-decoder for final object prediction. This design enables efficient and robust multispectral detection under challenging conditions.
Figure 3: Illustration of key components in the proposed framework. (a) Multi-Scale Feature Distillation (MSFD) module that enhances RGB features using multi-branch depthwise convolutions, channel shuffle, and grouped projection. (b) Cross-Modal Channel Shuffle Gate (CCSG) for gated channel-wise interaction. (c) Cross-Modal Large Kernel Gate (CLKG) for capturing long-range structural features.
2.2 Multi-Scale Feature Distillation
@@
To encourage cross-group communication and improve information flow, we apply a channel shuffle operation:
$\tilde{\mathbf{F}}=\mathcal{S}(\mathbf{F}_{\text{cat}})$ (6)
where $\mathcal{S}(\cdot)$ indicates the channel shuffle operation.
Finally, a grouped $1\times 1$ convolution is used to fuse the multi-scale features across channels, yielding the distilled RGB representation:
$\hat{\mathbf{F}}_{\text{RGB}}=\mathcal{G}_{1\times 1}(\tilde{\mathbf{F}})$ (7)
where $\mathcal{G}_{1\times 1}(\cdot)$ represents a grouped point-wise convolution.
This process enables the MSFD module to encode both fine-grained and large-scale context. The enriched RGB representation $\hat{\mathbf{F}}_{\text{RGB}}$ is then passed to the subsequent fusion stage.
Table 2: Performance comparison on FLIR-aligned and LLVIP datasets. We report mAP@50, mAP@75, and overall mAP. Our method achieves the best performance on LLVIP dataset.
@@
2.3 Cross-Modal Channel Shuffle Gate
@@
where $\odot$ denotes element-wise multiplication, and $\sigma(\cdot)$ is a GELU [14] activation function, which provides smooth and differentiable gating. This allows the model to softly emphasize informative features through cross-modal channel interactions while suppressing redundant and noisy components.
@@
2.4 Cross-Modal Large Kernel Gate
@@
3 Experiment
3.1 Experimental Setup
@@
Table 3: Ablation study on the impact of backbone configurations for RGB and IR modalities on the FLIR-aligned dataset. We report accuracy and model complexity.
@@
4 Experimental Results and Analysis
As shown in Table 2, our method achieves competitive performance on FLIR-aligned and achieves a state-of-the-art performance on LLVIP. Despite using a lightweight RGB backbone, IC-Fusion consistently outperforms prior fusion methods, demonstrating the effectiveness of our approach in challenging visual conditions. To understand the contribution of each modality, we perform an ablation study with different RGB-IR backbone configurations. As presented in Table 3, using a stronger backbone for the IR modality (ResNet-50) and a lightweight one for RGB (ResNet-18) yields the best performance-to-complexity trade-off. In contrast, assigning a stronger backbone to the RGB modality leads to a notable performance drop, highlighting the importance of focusing on IR features. Interestingly, our asymmetric setup even outperforms the dual ResNet-50 configuration despite having fewer parameters and reduced computational cost. This result further validates the effectiveness of our IR-centric design. We further analyze model complexity in Table 4, comparing IC-Fusion against RT-DETR baselines. Multiply-accumulate operations (MACs) are computed using the fvcore library. Qualitative results on the FLIR dataset are visualized in Fig. 4. Compared to unimodal RGB and IR detectors, IC-Fusion produces more accurate and stable predictions in low-light scenes, especially for small or occluded objects. This demonstrates the model’s ability to effectively leverage modality complementarity for robust detection under real-world conditions.
Table 4: Comparison of Parameters and MACs with baseline.
@@ 4 Experimental Results and Analysis
5 Conclusion
In this letter, we proposed IC-Fusion, a multispectral object detection framework that leverages IR-centric feature fusion. To this end, we designed a lightweight yet effective fusion module that integrates complementary cross-modal features. Specifically, the module incorporates a Multi-Scale Feature Distillation (MSFD) block to enhance RGB features, followed by a cascade of Cross-Modal Channel Shuffle Gate (CCSG) and Cross-Modal Large Kernel Gate (CLKG) blocks, which emphasize informative features from each modality while suppressing redundant information through gated cross-modal interaction. Extensive experiments on multispectral detection benchmarks validate the effectiveness of our design.
Acknowledgment
This work was supported by the IEEE. The authors would like to thank … Note that the Acknowledgment section of IEEE Sensors Letters is rendered in scriptsize.

### v2

# Multispectral Detection Transformer with Infrared-Centric Feature Fusion

@@ Abstract
Multispectral object detection aims to leverage complementary information from visible (RGB) and infrared (IR) modalities to enable robust performance under diverse environmental conditions. Our key insight, derived from wavelet analysis and empirical observations, is that IR images contain structurally rich high-frequency information critical for object detection, making an infrared-centric approach highly effective. To capitalize on this finding, we propose Infrared-Centric Fusion (IC-Fusion), a lightweight and modality-aware sensor fusion method that prioritizes infrared features while effectively integrating complementary RGB semantic context. IC-Fusion adopts a compact RGB backbone and designs a novel fusion module comprising a Multi-Scale Feature Distillation (MSFD) block to enhance RGB features and a three-stage fusion block with a Cross-Modal Channel Shuffle Gate (CCSG), a Cross-Modal Large Kernel Gate (CLKG), and a Channel Shuffle Projection (CSP) to facilitate effective cross-modal interaction. Experiments on the FLIR and LLVIP benchmarks demonstrate the superior effectiveness and efficiency of our IR-centric fusion strategy, further validating its benefits. Our code is available at https://github.com/smin-hwang/IC-Fusion.
Index Terms:
@@ I Introduction
Despite the potential of multispectral object detection [1, 2, 3, 4, 5], effectively integrating the two modalities remains challenging due to their inherent spectral differences. To better understand the contribution of each modality, we perform a wavelet decomposition on RGB and IR images, as shown in Fig. 1. The RGB image primarily contains low-frequency semantic structures in the LL band. In contrast, the IR modality presents distinct object boundaries and contours, especially in the LH and HL sub-bands, which are crucial for robust localization [6] and shape understanding [7]. This observation highlights that IR images inherently encode sharper structural cues that are highly beneficial for object detection, particularly in edge and boundary regions.
@@
The effectiveness of IR features is further quantitatively validated in Table I, where unimodal IR detectors consistently outperform their RGB counterparts on the FLIR-aligned [8] benchmark across various backbones. For example, RT-DETR [9] trained solely on IR data achieves a mean Average Precision (mAP) of 43.6, significantly surpassing its RGB equivalent (33.5 mAP). These findings imply that heavy feature extraction from the RGB modality may be redundant and detection performance can be retained or improved, even with significantly reduced computation by relying on IR-centric representations and adopting a lightweight RGB backbone.
Motivated by these insights, we propose a multispectral object detection transformer, IC-Fusion, which incorporates a novel cross-modal fusion module to efficiently fuse information across modalities. To reduce redundancy and improve efficiency, we adopt a lightweight backbone for the RGB modality while preserving a deeper IR representation to retain structural details. Our design emphasizes the dominant structural contributions of the IR modality while enhancing the complementary semantic information from RGB. Specifically, we introduce a Multi-Scale Feature Distillation (MSFD) module to enrich RGB features with multi-scale spatial context, and adopt a three-stage fusion block consisting of a Cross-Modal Channel Shuffle Gate (CCSG), a Cross-Modal Large Kernel Gate (CLKG), and a Channel Shuffle Projection (CSP) module to effectively extract and fuse complementary cross-modal representations. Experiments on benchmark datasets validate the effectiveness of our design and further highlight the advantage of IR-centric fusion for efficient multispectral object detection.
TABLE I: Comparisons on the FLIR-aligned dataset. We report mAP@50, mAP@75, and mAP under IoU thresholds.
@@
Figure 2: Overview of (a) the proposed IC-Fusion framework and (b) the proposed cross-modal feature fusion module composed of CCSG, CLKG, and CSP modules.
II Proposed Method
II-A Overview
We propose IC-Fusion, a multispectral object detection transformer framework that effectively integrates RGB and IR features through structured cross-modal fusion. As shown in Fig. 2, the framework consists of three main components: dual-stream backbones, a fusion module, and a DETR-based transformer encoder-decoder [9, 12, 13, 14].
@@ II-A Overview
where $\mathcal{F}(\cdot)$ represents the cross-modal fusion process, which sequentially applies a Cross-Modal Channel Shuffle Gate (CCSG) for feature interaction, a Cross-Modal Large Kernel Gate (CLKG) for long-range contextual alignment, and a Channel Shuffle Projection (CSP) module. The output of this three-stage fusion block then undergoes a channel split into two parallel feature streams, which are subsequently element-wise added to form the final fused feature map. This design facilitates the integration of cross-modal information, thereby enhancing feature richness and training stability by enabling the model to selectively propagate refined features.
The fused output integrates complementary semantics from both modalities while suppressing redundant information, and is subsequently passed to an efficient encoder [9], followed by a DETR-decoder for final object prediction. This design enables efficient and robust multispectral detection under challenging conditions.
Figure 3: Illustration of key components in the proposed framework. (a) Multi-Scale Feature Distillation (MSFD) module that enhances RGB features using multi-branch depthwise convolutions, channel shuffle, and grouped projection. (b) Cross-Modal Channel Shuffle Gate (CCSG) for gated channel-wise interaction. (c) Cross-Modal Large Kernel Gate (CLKG) for capturing long-range structural features. (d) Channel Shuffle Projection (CSP) module for efficient feature projection and distillation.
II-B Multi-Scale Feature Distillation
@@
TABLE II: Performance comparison on FLIR-aligned and LLVIP datasets. We report mAP@50, mAP@75, and overall mAP. Our method achieves the best performance on LLVIP dataset.
@@
Finally, the concatenated features $\mathbf{F}_{\text{cat}}$ are passed to a Channel Shuffle Projection (CSP) module for final feature distillation. As illustrated in Fig. 3(d), the CSP module is designed to efficiently fuse features across channels. It first applies a channel shuffle operation to enable effective information flow across different channel groups. Subsequently, it employs two sequential grouped $1\times 1$ convolutions and GELU [18] activation to project and distill the features in a computationally lightweight yet expressive manner. The operation of CSP can be formalized as:
$\hat{\mathbf{F}}_{\text{RGB}}=\mathcal{G}_{2}\left(\sigma\left(\mathcal{G}_{1}\left(\mathcal{S}(\mathbf{F}_{\text{cat}})\right)\right)\right)$ (6)
where $\mathcal{S}(\cdot)$ denotes the channel shuffle operation, $\sigma(\cdot)$ is the GELU activation function, and $\mathcal{G}_{1}(\cdot)$ and $\mathcal{G}_{2}(\cdot)$ represent the first and second grouped $1\times 1$ convolutions, respectively. This composite process yields the distilled RGB representation.
This process enables the MSFD module to encode both fine-grained and large-scale context through its multi-branch structure, with the final distillation handled by the CSP module. The enriched RGB representation $\hat{\mathbf{F}}_{\text{RGB}}$ is then passed to the subsequent fusion stage.
II-C Cross-Modal Channel Shuffle Gate
@@
where $\odot$ denotes element-wise multiplication, and $\sigma(\cdot)$ is a GELU, which provides smooth and differentiable gating. This allows the model to softly emphasize informative features through cross-modal channel interactions while suppressing redundant and noisy components.
@@
II-D Cross-Modal Large Kernel Gate
@@
III Experiments
III-A Experimental Setup
@@
TABLE III: Ablation study on the impact of backbone configurations for RGB and IR modalities on the FLIR-aligned dataset.
@@
TABLE IV: Ablation study on the contribution of each proposed component. Results are reported on the FLIR-aligned dataset.
MSFD CCSG CLKG mAP50 mAP75 mAP
80.7 40.2 44.1
81.0 41.3 44.7
82.4 42.3 45.4
83.9 42.6 45.8
III-B Experimental Results and Analysis
As shown in Table II, our method achieves competitive performance on FLIR-aligned and achieves a state-of-the-art performance on LLVIP. Despite using a lightweight RGB backbone, IC-Fusion consistently outperforms prior fusion methods, demonstrating the effectiveness of our approach in challenging visual conditions. To understand the contribution of each modality, we perform an ablation study with different RGB-IR backbone configurations. As presented in Table III, using a stronger backbone for the IR modality (ResNet-50) and a lightweight one for RGB (ResNet-18) yields the best performance-to-complexity trade-off. In contrast, assigning a stronger backbone to the RGB modality leads to a notable performance drop, highlighting the importance of focusing on IR features. Interestingly, our asymmetric setup even outperforms the dual ResNet-50 configuration despite having fewer parameters and reduced computational cost. This result further validates the effectiveness of our IR-centric design.
Table IV presents a detailed ablation study on the contribution of each proposed component. The results clearly show that each module incrementally improves performance, validating their individual effectiveness and necessity within our fusion framework. We further analyze model complexity in Table V, comparing IC-Fusion against RT-DETR baselines. Multiply-accumulate operations (MACs) are computed using the fvcore [23] library. Qualitative results on the FLIR dataset are visualized in Fig. 4. Compared to unimodal RGB and IR detectors, IC-Fusion produces more accurate and stable predictions in low-light scenes, especially for small or occluded objects. This demonstrates the model’s ability to effectively leverage modality complementarity for robust detection under real-world conditions.
TABLE V: Comparison of Parameters and MACs with baseline.
@@ III-B Experimental Results and Analysis
IV Conclusion
In this letter, we proposed IC-Fusion, a multispectral object detection framework that leverages IR-centric feature fusion. To this end, we designed a lightweight yet effective fusion module that integrates complementary cross-modal features. Specifically, the module incorporates a Multi-Scale Feature Distillation (MSFD) block to enhance RGB features, followed by a cascade of a Cross-Modal Channel Shuffle Gate (CCSG), a Cross-Modal Large Kernel Gate (CLKG), and a Channel Shuffle Projection (CSP) module, which emphasize informative features from each modality while suppressing redundant information through gated cross-modal interaction. Extensive experiments on multispectral detection benchmarks validate the effectiveness of our design.